Direct Labor Variances

La Barte Company manufactures commuter bicycles from recycled materials. The following data for July of the current year are available:

Quantity of direct labor used 690 hrs.
Actual rate for direct labor $12.40 per hr.
Bicycles completed in July 320 bicycles
Standard direct labor per bicycle 2 hrs.
Standard rate for direct labor $12.60 per hr.

a. Determine the direct labor rate variance, time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct labor rate variance $
Direct labor time variance
Total direct labor cost variance $

b. How much direct labor should be debited to Work in Process?
Answer #1

Labor rate variance = (SR-AR) *AH

= (12.60-12.40)*690

= -138 Favourable (minus)

Labor time variance = (SH-AH) *SR

= (320*2 - 690)*12.60

= 630 Unfavorable

Labor cost variance = 138 F + 630 U

= 492 Unfavorable

Direct labor debited to work in process

= 320*2*12.60

= 8064